OpenStudy (anonymous):

Square root of (x+deltaX)+ 2 - square root of (x-2)/ delta x

OpenStudy (john_es):

What is the question?

OpenStudy (anonymous):

Find the lim as delta x approaches 0

OpenStudy (john_es):

Do you mean \[\sqrt{x+\Delta x}+2-\frac{\sqrt{x-2}}{\Delta x}\]

OpenStudy (john_es):

\[\frac{\sqrt{x+\Delta x}+2-\sqrt{x+2}}{\Delta x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but for the first sqrt its (x+deltax)+2

OpenStudy (john_es):

Ok this is an indetermination of the type 0/0. You can use this method, \[\frac{\sqrt{x+\Delta x+2}-\sqrt{x+2}}{\Delta x}\cdot\frac{\sqrt{x+\Delta x+2}+\sqrt{x+2}}{\sqrt{x+\Delta x+2}+\sqrt{x+2}}\]

OpenStudy (anonymous):

the answer is apparently \[\frac{ 1 }{ 2\sqrt{x+2}}\]

OpenStudy (anonymous):

alright ill try it

OpenStudy (john_es):

Once you multiply the factors, you'll obtain, \[\frac{\Delta x}{\Delta x(\sqrt{x+\Delta x+2}+\sqrt{x+2})}=\frac{1}{(\sqrt{x+\Delta x+2}+\sqrt{x+2})}\]And yes, now take the limit Delta x tends to 0 and you'll obtain the answer.

OpenStudy (john_es):

;)

OpenStudy (anonymous):

yup it worked perfectly

OpenStudy (john_es):

Use this "trick" when you have limits of the same type (also if you see derivatives, you'll see that it is possible use the L'Hopital rule)

OpenStudy (anonymous):

we havent gotten to derivatives in ap calc yet so yeah... ahah